Question 1 A professional organizer of children’s summer camp luggage gets paid an average of $320 per hour with a standard deviation of $41.75 per hour. How much would an organizer in the top 13% of pays receive per hour for their services? Assume the population is normally distributed. Round to the hundredths place, as needed. Use the tables in the book and not an outside source. Please show work for the opportunity to earn partial credit on this question.
Answer #1

Given that,

mean = = $320

standard deviation = =$41.75

Using standard normal table,

P(Z > z) =13 %

= 1 - P(Z < z) = 0.13

= P(Z < z ) = 1 - 0.13

= P(Z < z ) = 0.87

= P(Z <1.13 ) = 0.87  

z = 1.13 (using standard normal (Z) table )

Using z-score formula  

x = z * +

x= 1.13*41.75+320

x= 367.1775

x=367.18